OpenStudy (anonymous):

What is the value of expression? 5−8 • 5−4 A. 532 B. 54 C. 5−4 D. 5−12

Nnesha (nnesha):

\[\huge\rm 5^{-8} \times 5^{-4}\] like this ?

OpenStudy (anonymous):

yes

Nnesha (nnesha):

when we multiply`same bases` we should `ADD` their exponents \[\huge\rm x^m \times x^n=x^{m+n}\]

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

every time i solve this i get 0.000000004

Nnesha (nnesha):

don't use a calculator.

OpenStudy (anonymous):

ok

Nnesha (nnesha):

bases are the same then add their exponents